OpenStudy (anonymous):

What is the surface area of a cone with a radius of twelve centimeters and a slant height of 8 centimeters? Use three point one four for pi.

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

let h(s) be the slant height. \[SA = \frac{ 1 }{ 2 }2\pi r h _{s}+\pi r ^{2}\]

OpenStudy (anonymous):

1/2 times circumference times slant height + area of base

OpenStudy (anonymous):

You can see that the 1/2 and 2 will cancel eachother out

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

AKA Do my homework for me. I gave you the equation, it is your job now to plug the data in and solve